OpenStudy (anonymous):

Does ln|1-(e^-x)| = -x-ln|(e^s)-1|? I graphed it but it does not look the same. The problem is integrate dx/((e^x)-1), I got -x-ln|(e^s-1)| + C as my answer, but the back of the book says it is ln|1-(e^-x)| + C.

OpenStudy (anonymous):

Where does the \(s\) come from? If you integrate a function of \(x\) with respect to \(x\), you can only get another function of \(x\). \[\int\frac{dx}{e^{x}-1}\] Let \(u=e^{x}-1~\Rightarrow~e^x=u+1\\ du=e^x~dx~\Rightarrow~dx=\dfrac{du}{u+1}\) \[\int\frac{\frac{du}{u+1}}{u}\\ \int\frac{du}{u(u+1)}\] Partial fraction decomposition (leaving the details to you): \[\frac{1}{u(u+1)}=\frac{1}{u}-\frac{1}{u+1}\] \[\int\frac{du}{u}-\int\frac{du}{u+1}\\ \ln|u|-\ln|u+1|+C\\ \ln|e^x-1|-\ln|e^x|+C\\ \ln|e^x-1|-x+C\] You seem to be off by a minus sign on the first term. As for why the book has that answer: \[\begin{align*}\ln|e^x-1|-\ln|e^x|+C&=\ln\left|\frac{e^x-1}{e^x}\right|+C\\ &=\ln\left|1-e^{-x}\right|+C\end{align*}\]

OpenStudy (anonymous):

The s was a typo, lol. I rewrote the numerator as e^s - e^s + 1 and didn't use partial fractions, but I think it was much easier to see it using PF. Thanks!